Weekly meeting on Xboxes
https://cern.zoom.us/j/6910405761
Xboxes weekly meeting. 10.01.2024
Attending: Paz Alonso Arias (PAA), Nuria Catalan Las Heras (NClH), Chiara Marrelli (CM), Matteo Volpi (MV), Carmelo Barbagallo (CB), Alexei Grudiev (AG), Anisullah Baig (AB), Ping Wang (PW), Igor Syratchev (IS), Olivier Brunner (OB)
Xbox2. PAA presents a summary on the Xbox status before going on Christmas break:
· Xbox2 restarted in december after Scandinova modulator and SSA unit where fixed
· Changed length of pulse to 50ns to start ramping in power with small pulses, then we will increase pulse width
· Installed flow meters at input of structure, and temperature sensors at input/output and surface of structures. Data analysis by Mareike showed unexpected results.
· During Christmas break there were some electrical cuts, and the solenoid temperature increased to 107K, we lost superconductivity (~1 week needed before start operating again)
NClH asks PAA to contact Microwave Amps to check if they have received the SSA units that were sent to repair.
PAA comments the cryocooler does not work, temperatures are not descending correctly. Could be a lack of helium liquid. PAA asks who to contact with to check helium. NClH: Torsten Koettig, if not Sergio Calatroni (cryogenics lab).
NClH asks about temperature sensors and results by Mareike. PAA comments on low accuracy of temperature sensors. We need to check that. PAA to double DC-Down channel and attenuate it with a resistor for future measurements.
Xbox3-HE klystrons. PAA presents summary of power measurements, including weird behaviour in frequency in comparison to Canon measurements. CM asks if Canon did perform measurements of saturation power. PAA and IS confirm that they did not (they fix power and move in frequency/voltage, we did both). IS explains that a good explanation for weird behaviour could be presence of harmonic signal, we need to check.
PAA comments on some changes to solve problems with vacuum, including change of load and HPDC (PAA characterised the Sparameters before installing). NCHL asks if we swapped klystrons, PAA confirms we didn’t. It’s decided that if we still have weird results after calorimetry, we will need to swap them to see if it’s a problem of the modulator.
NCHL asks if we have everything for doing calorimetry. PAA says we are still waiting for material for calorimetry, temperature sensors and flow meters. We already have thermal jackets and controller to isolate the loads and warm them to control temperature. There is a discussion between NClH, AG and IS about whether it’s necessary to warm the loads and control temperature. It’s decided to not plug the thermal jackets and just isolate the loads, as AG says we could have some vacuum leaks if we warm them. In case, we can always warm them up later.
Others.
NClH informs an intern is coming on February, to continue data analysis started by Mareike, to be supervised by PAA.
NClH asks about status of BOC. AG and PW confirm it is ready, but they want to do one more metrology measurement to confirm what kind of mechanical errors degrade performance. They won’t modify anything, so it can be installed as soon as we want.
There is a small discussion on where to install the BOC: xbox2 or xboc3. AG thinks XBOX2 would be better as 50MW are needed (as high as possible), but we had previously discussed to install it in Xbox3 as DUT or as pulse compressor. IS says that the correction cavities in Xbox2 are right now tuned to the current pulse compressor, so we would need to change that to install the BOC on XBOX2. PW will do some analysis on compressed pulse shapes to evaluate where is better, before a final decision is made and we do integration.
Melbourn news. MV says next week they can start conditioning of line without the structure. They have calibrated the HPDC, and confirmed that they don’t have leaks in the load. They expect problems with pulse compressor.
Radiation monitors have arrived and are installed. Interlock crate still pending. MV says they also have a summer student, that is supposed to apply to come to CERN during summer. NCLH asks what they want to do, MV says she knows about Xboxes, controlling, VNA measurements.
Dielectric measurements. NCLH comments dielectric machine has been installed by PAA and Pedro. She asks if Adnan could do measurements of dielectrics? AG says Adnan is doing very different things now, but maybe either him or PW in next months could do measurements. IS says AB could do also measurements, as they have their own material to test. AB comments they still need to get the sample. NClH says to ask Pedro about optimal thickness of sample to test.
PAA says she has made a simple procedure to follow for measurements. The machine works ok, but measurements are a bit noisy. Pedro is working on how to improve setup to fix machine, VNA and cables and reduce vibrations.
OB asks about news of brazing tests. NClH says that there are some results presented by Pedro on CLIC mini week and to ask him. Summary: both brazing materials were ok. But preferred option will be gold.
MV questions about UPS installed to avoid destroying the SSA units when there is a power cut. NClH confirms it was Alan to do that work (he purchased an ondulator sized for the power we have), and to ask him.